OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 2) (5, 5)

OpenStudy (***[isuru]***):

first find the slope.. can u find it ? \[slope = \frac{ y_{2} - y_{1} }{ x_{2} -x_{1} }\]

OpenStudy (***[isuru]***):

then u can use point slope form... since u know the coordinates..

OpenStudy (***[isuru]***):

\[slope = \frac{ 5-2 }{ 5-6 }\] \[slope =-3\]

OpenStudy (***[isuru]***):

now the standered form of a function is y = mx + b where m = slope b = intercept... now since u know the slope y= -3x + b to find be u can use (6 , 2) if u plug them in to ur equation 6 =-3(2) + b b = 12 so... the equation is y = -3x + 12 got it :) ?

OpenStudy (anonymous):

thank you